The probability of rolling a number greater than 5 on one roll of a die?
Only 6 is greater than 5 on a standard cubic die. The chance of rolling a 6 is one in six.
Two dice are thrown what are the odds in favor of rolling a score of six?
The first die can come up in any one of 6 ways.
The second die can come up in any one of 6 ways.
There are (6 x 6) = 36 different ways that two dice can fall.
There are 5 ways that they can come up as a six:
1 + 5
5 + 1
2 + 4
4 + 2
3 + 3
The probability of rolling a 6 is (5/36) = 13.89 %
The odds in favor are 5 to 31.
1.96
Tackling multiple choice questions relies primarily on the memory process called?
Tackling multiple choice questions relies primarily on the memory process called recognizing.
George Wickham did not attend the ball at Netherfield, because he did not want to meet Mr. Darcy
From a deck of 52 cards one card is drawn what is the probability that the card will be a red ten?
Since there are 2 red tens in a 52 card deck, probability of drawing one of the 2 red tens is 2/52 or 1/26 or 0.0385.
Independent. The chance the die are either a total of 11 or one will be a 4 are not related.
---------------------------------------------------------------------------------------------
There are only two ways the sum of the resulting numbers of two dice thrown give 11: (5,6) and (6,5).
The probability for event A is: P(A) = 2(1/6)(1/6) = 1/18 = 0.05555... ≈ 5.56%
Given event A, there is no possibility that one of the two dice show a 4.
The conditional probability is, P(B|A) = 0.
So, event B is dependent of event A.
1
85/500 = 17%
What is your chance of getting a diamond out of 52 cards?
There is one chance in four of being dealt a diamond from a 52 card deck.
If John tosses four fair coins what is the probability that all four coins will land tails up?
I don't know! I asked you!
What is the probability of a coin landing on its side?
Prob. of Landing on one side = 0.5
Prob. of Landing on the other = 0.5
Therefore, Prob. of Landing on one side or the other = 0.5 +0.5 = 1
:)
I think you mean 'landing on its edge'. In that case it depends on the coin (some even have rounded edges).
Interestingly what face a coin lands on is not exactly 50/50. Each person has an individual flipping-style which leads to a different side being preferred (not explicitly heads vs. tails but in 'landing with the same face upwards as was upwards at the beginning of the flip' vs. 'landing with the other face upwards to the way it started'). On average (with most people) the coin lands the same way up as it started more often than not (you'll have to do hundreds and hundreds of flips before you notice this).
If (like me) you want to be really pedantic you would also take into account the very slight possibility that the coin won'tland at all.
What is the equation of probability?
The probability of an event occurring can be found by dividing the number of favorable outcomes (what you want to happen) by the number of possible outcomes
number of favorable outcomes
probability = _________________________
number of possible outcomes
What is the probability of drawing a queen or a club from a deck of cards?
Probability (P) of A or B is: P(A) + P(B) - P(A and B). Apply to the question is: P(Q) + P(Club) - P (Q&Club) = 4/52 + 13/52 - 1/52 = 16/52 = 4/13 or .3077 or 30.77%.
At Z = 0.25, the area (larger) to the left of the line is 0.5987. The smaller area to the right of the line is 1 - 0.5987 = 0.4013. Therefore, the smaller section corresponds to 40.13% of the area.
From the table in the related link, Z = 1.43, area to left is 0.9236. This is read directly from the table.
What is the Difference between z scores and percentile ranks?
First let's define both, that will help to see the difference.
1.A percentile is a measure that lets us know what percent of the total frequency scored below that measure. A percentile rank is the percentage of scores that fall below a given score. Here is how that works.
Given a score, call it S and a total of n scores we are looking at, we find the number of scores below S and divide that by n. Next multiply that by 100 and you have the percentile rank.
Now a z score is the number of standard deviations from the mean.
Say the mean is M and your score is S as above. Let sigma be the standard deviation of the distribution. Then z=(S-M)/sigma.
So let's say the mean M is 100 and sigma is 15. S is 132, you did better than average!
So z=(132-100)/15=2.13
If 60 percent of the people scored less than you, then you are in the 60th percentile.
Furthermore, lets say, there were 100 people taking test, then 60 of them scored less than you. Your percentile ranking is (60/100)x100=60
So both are measures of where your results falls in a distribution. z scores are often used for probability of a certain result. Percentile ranks are often used in looking at standardized test results or growth data. One can convert from one to the other.
I have given a conversion table link below
http://www.acposb.on.ca/conversion.htm
prob of odd number: 6/12
prob less than 4: 3/12
prob odd AND less than 4: 2/12
18
What is The probability of being born on 1st of January?
Births are not distributed randomly across the year: there are seasonal variations as well as differences between days of the week. Where maternity staff like to celebrate New Year's Eve, there may be fewer people on duty on January 1 and so there may well be fewer births. Unfortunately, comprehensive information on such variations is not available.
If you assume that each day of the year is equally likely, the probability is 1/365. Adjusting for leap years changes that to 1/365.25. Further adjusting for Centuries gives 1/365.2425
Prob(The one block is blue) = number of blue blocks/total number of blocks = 8/25 = 32%
The probability that the birthdays of five persons chosen at random will fall in twelve different calender months is zero.
You would need at least twelve persons to have a non zero probability.
How many 4 digit EVEN numbers have all 4 digits distinct?
4 digit number having 0 at last position = 9*8*7
=504
4 digit even number having 2 or 4 or 6 or 8 at last position = 8*8*7*4
=1792
(in first position, 0 cannot be fix and in last position one of digit 2,4,6,8 will be fix,
so, number of choice left for first position is 8)
So, total 4 digit even number possible = 504 + 1792
= 2296